OpenStudy (anonymous):

cos(x/3)cos(x/3)=(1/2)(1+cos(2/3)) true or false? i really have no idea!!!

zepdrix (zepdrix):

Did you mean to write this?\[\large \cos\left(\frac{x}{3}\right)\cos\left(\frac{x}{3}\right)=\frac{1}{2}\left[1+\cos\left(\frac{2x}{3}\right)\right]\] There is no x on the right side of your equation, so I wanted to see if this is what you meant to write.

OpenStudy (anonymous):

Yes! thankyou

zepdrix (zepdrix):

Here is our `Half Angle Formula for Cosine`:\[\large \cos^2 \theta=\frac{1}{2}\left[1+\cos2\theta\right]\]A good one to remember. Hmmmmmm, does it look like the rule is being followed in this case? Think about it a sec! :)

OpenStudy (anonymous):

yeah...haha

zepdrix (zepdrix):

Yah it looks like it's true doesn't it?? :O Our angle is x/3 in this case. It appears to be twice that angle on the right.. I think we have a winner!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

give him a medal atleast geez.